Essence of Skunk Fragrances, Ltd., sells 7,400 units of its perfume collection each year at a price per unit of $690. All sales are on credit with terms of 1/15, net 60. The discount is taken by 55 percent of the customers.

Calculate the average collection period. (Round your answer to 2 decimal places, e.g., 32.16.)

Average collection period              days

What is the receivables turnover? (Use 365 days a year. Do not round intermediate calculations and round your answer to 3 decimal places, e.g., 32.161.)

Receivables turnover              times

What is the amount of the company’s average accounts receivable? (Do not round intermediate calculations and round your answer to 2 decimal places, e.g., 32.16.)

Average accounts receivable            $
Answer #1
Average collection period
Average collection period = 15*55% + 60*45% = 35.25
                     35.25 Days
Receivables turnover
Receivables turnover = 365/35.25 = 10.355
                   10.355 Times
Average accounts receivable
Average collection period = Average Accounts receivable/ (Annual Credit Sale/365)
35.25= Average Accounts receivable / (5106000/365)
Average Accounts receivable = 13989.04*35.25 493113.66
Average accounts receivable $              4,93,113.66

--

Hope you Understood.
If you have any doubt please leave a comment.

Thank you.